OpenStudy (anonymous):

How do I parametrize the curve of intersection between the surfaces z=x^2+y^2 and 2x-4y-z-1=0?

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (nikvist):

\[z=x^2+y^2\quad;\quad 2x-4y-z-1=0\quad,\quad z=2x-4y-1\] \[2x-4y-1=x^2+y^2\] \[-1=x^2+y^2-2x+4y=(x-1)^2+(y+2)^2-1-4\] \[\Rightarrow (x-1)^2+(y+2)^2=4\quad\Rightarrow x=1+2\cos{t}\,,\,y=-2+2\sin{t}\] \[z=(1+2\cos{t})^2+(-2+2\sin{t})^2=9+4\cos{t}-8\sin{t}\,,\,t\in[0,2\pi]\]

OpenStudy (anonymous):

Thank You!